### v1

# Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity

[…]
Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity
Genki Shimizu, Taro Toyoizumi
Short-term synaptic plasticity (STP) is traditionally viewed as a purely presynaptic mechanism that filters incoming spike trains without detecting postsynaptic activity. Recent experiments, however, have revealed an associative form of STP in which presynaptic release probability is modified in concert with long-term potentiation, suggesting a richer computational role for presynaptic plasticity. Here we develop a normative theory of such associative STP using an information-theoretic framework. Extending Fisher-information-based learning rules to synapses with Tsodyks–Markram dynamics, we derive analytic update rules for both baseline synaptic strength and release probability that maximize encoded stimulus information under resource constraints. The resulting learning rules decompose into a conventional postsynaptic term that tracks local firing and a distinct presynaptic term. The latter term exhibits a novel phase-advanced structure relative to firing rate and selectively detects stimulus onset; critically, it is within this presynaptic component that the plasticity of baseline strength and release probability differ. When stimulus drive varies on timescales slower than the EPSP time constant, this onset sensitivity biases optimal connectivity toward anti-causal associations, strengthening synapses from neurons activated later in a stimulus to those activated earlier. In recurrent circuits, networks optimized under these rules develop ramp-like representations during sustained stimulation and support reverse replay sequences when external drive is removed. Linear-response analysis further shows that STP endows presynaptic drive with frequency-dependent phase selectivity and that constraints on total release probability systematically modulate the strength of temporal asymmetry. Together, these results provide a principled account of the computational role of associative STP and highlight presynaptic plasticity of release probability as a key substrate for rapidly reconfigurable temporal coding.
1 Introduction
Learning through experience is mediated by synaptic plasticity, the activity-dependent modification of connection strengths between neurons. Classical long-term potentiation (LTP) has been extensively studied as the primary mechanism for associative learning, where correlated pre- and postsynaptic activity leads to persistent strengthening of synapses through increased postsynaptic receptor expression [bliss1973Longlasting, bi1998Synaptic]. This Hebbian plasticity enables the formation of neural assemblies that encode relationships between concepts. However, LTP typically requires tens of minutes to hours for full expression and once expressed, can persist from hours to days or longer, which cannot fully account for flexible behavioral adaptation occurring on faster timescales.
[…]
To address this question, we adopt an information-theoretic framework. Neural populations have been shown to optimize their representations for efficient information transmission [linsker1988Selforganization], and Hebbian-like plasticity rules can emerge from that optimization process [linsker1988Selforganization, toyoizumi2005Generalized, toyoizumi2005Spiketiming, toyoizumi2006Fisher]. While STP has been shown to enhance information processing in various contexts [rotman2011Shortterm, regehr2012ShortTerm], the impact of associatively learning STP parameters on neural encoding remains unexplored. In this article, we focus on the functional significance of associative STP from the perspective of efficient neural representations. In particular, we use Fisher information as a local and tractable measure of encoding efficiency [dayan2001Theoretical, brunel1998Mutual].
To this end, we extend the Fisher information optimization framework [seung1993Simple, toyoizumi2006Fisher] from static synaptic weights to dynamic synapses characterized by activity-dependent vesicle release. Specifically, we adopt the Tsodyks-Markram model [tsodyks1997Neural, tsodyks1998Neural] of short-term synaptic dynamics and derive learning rules for both baseline synaptic weight and release probability. We find that this optimization leads to temporal asymmetry: anti-causal connections (where presynaptic neurons lag behind postsynaptic neurons) are preferentially strengthened, contrasting sharply with classical STDP. When stimulus drive varies on timescales slower than the EPSP time constant, the resulting networks emphasize stimulus offsets through “ramp-up” representations and naturally support “reversal” connections of experienced sequences. Furthermore, the extent of temporal asymmetry depends on constraints on release probability, potentially explaining state-dependent differences in replay directionality observed during wakefulness versus sleep. Our framework provides a principled understanding of how associative short-term plasticity shapes neural representations and suggests its previously unappreciated role in rapid learning and memory consolidation.
2 Model
2.1 Neuron Model
Following the setting of [seung1993Simple, toyoizumi2006Fisher], we consider a network of stochastically firing spiking neurons.
The membrane potential $u_{i}(t)$ of each neuron $i$ is determined by the sum of stimulus-dependent external input $h_{i}(t,\theta)$ and recurrent synaptic inputs from other neurons:
$u_{i}(t)=h_{i}(t,\theta)+\sum_{j=1}^{N}\sum_{t_{j}^{f}}\epsilon(t-t_{j}^{f})w_{ij}(t_{j}^{f})$ (1)
where $w_{ij}(t)$ represents the time-dependent synaptic strength from neuron $j$ to neuron $i$ at time $t$, more precisely, the amplitude of the excitatory postsynaptic potential (EPSP) jump that would occur if a spike arrived at time $t$. The function $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$ represents the causal EPSP kernel with membrane time constant $\tau_{m}$, where $\Theta(t)$ is the Heaviside step function.
For stochastic firing, we assume that the instantaneous firing rate $\rho_{i}(t)$ is determined by a nonlinear function $g(u)$ of the membrane potential, without considering refractory periods or membrane potential reset: $\rho_{i}(t)=g(u_{i}(t))$. Consequently, neuronal firing follows an inhomogeneous Poisson process. We primarily consider an exponential activation function $g(u)=g_{c}e^{\beta(u-u_{c})}$, where $\beta$ controls the gain, and $u_{c}$ is the threshold. Our results remain qualitatively similar for other reasonable activation functions, such as the sigmoid function $g(u)=g_{M}[1+e^{-\beta(u-u_{c})}]^{-1}$. This model can be viewed as a special case of the spike response model. Furthermore, we assume that spike generation is conditionally independent across neurons given the membrane potential and spike history, thereby neglecting common noise sources.
2.2 Synapses with Short-Term Dynamics
Synaptic strengths $w_{ij}(t)$ follow Tsodyks-Markram short-term plasticity dynamics [tsodyks1997Neural, tsodyks1998Neural]. In this model, the synaptic efficacy is expressed as the product of a facilitation factor $u_{ij}(t)$ and a depression factor $d_{ij}(t)$:
$w_{ij}(t)=w_{ij}^{0}u_{ij}(t)d_{ij}(t)$ (2)
where $w_{ij}^{0}$ denotes the baseline synaptic strength, $u_{ij}(t)$ corresponds to the vesicle release probability, and $d_{ij}(t)$ represents the fraction of available neurotransmitter resources. Variables $u_{ij},d_{ij}$ are normalized such that $u_{ij},d_{ij}\in[0,1]$.
For simplicity, we assume that the facilitation time constant $\tau_{f}$ is sufficiently short compared to the timescales of interest, allowing us to treat the facilitation variable as constant during stimulus presentation episodes: $u_{ij}(t)\equiv U_{ij}$ (the validity of this assumption is discussed later in the Discussion section; see Sec. 4.4).
Under this assumption, the dynamics of synaptic strength are governed solely by the evolution of the depression variable:
$\dot{d}_{ij}=\frac{1-d_{ij}}{\tau_{d}}-U_{ij}d_{ij}(t^{-})\delta(t-t^{\text{spike}})$ (3)
where $d_{ij}(t^{-})$ denotes the value immediately before time $t$ (unaffected by spike input at time $t$), and $\delta(t)$ is the Dirac delta function representing presynaptic spikes. The effective synaptic strength $w_{ij}(t)=w_{ij}^{0}U_{ij}d_{ij}(t)$ then follows:
$\dot{w}_{ij}=\frac{w_{ij}^{0}U_{ij}-w_{ij}(t)}{\tau_{d}}-U_{ij}w_{ij}(t^{-})\delta(t-t^{\text{spike}})$ (4)
In this formulation, the baseline strength $w_{ij}^{0}$ characterizes the postsynaptic component (e.g., receptor expression level), while the release probability $U_{ij}$ governs the presynaptic dynamics of neurotransmitter release.
2.3 Fisher Information
We consider a population of neurons receiving parameter-dependent input $h_{i}(t,\theta)$, where $\theta$ is the stimulus parameter to be encoded. $x_{i}(t)=\sum_{f_{i}=1}^{n_{i}}\delta(t-t_{i}^{f_{i}})$ denotes the output spike train of neuron $i$ and the complete spike histroy is denoted by $X(t)=\{x_{i}(t^{\prime})\mid i=1,\ldots,N;\,0\leq t^{\prime}\leq t\}$.
As calculated in [toyoizumi2006Fisher], the Fisher information, which quantifies how accurately a downstream decoder can estimate $\theta$ from the population activity, is given by:
$J$ $=-\left\langle\frac{\partial^{2}\log P(X)(T)}{\partial\theta^{2}}\right\rangle_{X(T)}$ (5)
$=\int_{0}^{T}dt\sum_{i=1}^{N}\left\langle\left[h^{\prime}_{i}(t)\frac{g^{\prime}_{i}(t)}{g_{i}(t)}\right]^{2}\rho_{i}(t)\right\rangle_{X(t)}$ (6)
where $h^{\prime}_{i}(t)=\partial h_{i}(t,\theta)/\partial\theta$ represents the sensitivity of the external input to the parameter, $g^{\prime}_{i}(t)=dg(u)/du|_{u=u_{i}(t)}$ is the derivative of the activation function evaluated at the current membrane potential, and $\langle\cdot\rangle_{X(t)}$ denotes the average over the stochastic spike history $X(t)$.
While [toyoizumi2006Fisher] derived learning rules for time-invariant synaptic weights that maximize Fisher information, our goal is to extend this framework to synapses with short-term dynamics. Specifically, we aim to derive learning rules for both the baseline synaptic strength $w_{ij}^{0}$ and the release probability $U_{ij}$ that maximize the Fisher information in networks with dynamic synapses. Although the initial postsynaptic response scales with the overall product of $w_{ij}^{0}U_{ij}$, these parameters play distinct roles: $U_{ij}$ additionally governs how synaptic efficacy evolves during presynaptic spike trains through STP dynamics, justifying their separate optimization.
3 Results
We now analyze how short-term synaptic plasticity shapes the optimal learning rule based on Fisher information. First, we obtain explicit gradient expressions for the parameters of the Tsodyks–Markram model and rewrite them in a Hebbian-like form that isolates the contributions of presynaptic dynamics. We then examine the resulting learning rule to clarify how the learning rule under short-term depression determines the temporal and frequency selectivity of synaptic modifications. Finally, we apply this framework to a ring network supporting traveling-wave activity and demonstrate how optimizing information transmission sculpts synaptic structure and sequential neural activity in a concrete circuit model.
3.1 Derivation of Learning Rules
In this subsection, we derive synaptic learning rules for short-term depressing synapses by differentiating the Fisher information with respect to the baseline synaptic strength $w_{ij}^{0}$ and the release probability $U_{ij}$. We first present an exact gradient formula that holds for arbitrary recurrent coupling, expressed in terms of a point-process score term and an eligibility trace that can be computed online from single network simulations. We then obtain a tractable analytic approximation in the weak-coupling regime ($w_{ij}(t)\ll 1$), where the gradient reduces to a correlation between a postsynaptic information-weighting factor and a filtered effective presynaptic drive. To isolate the role of short-term depression, we introduce normalized sensitivity functions whose dynamics are closed under weak coupling, yielding distinct presynaptic components for $w_{0}$ and $U$. Finally, we interpret the resulting learning rules in a “Hebbian” factorized form and analyze how the presynaptic term becomes onset-sensitive (phase-advanced) under step and oscillatory inputs, providing intuition for the anti-causal bias in the connectivity patterns implied by Fisher-information optimization.
3.1.1 Exact gradient for arbitrary recurrent coupling
We first state an exact expression for the gradient of the Fisher information that holds for arbitrary recurrent coupling strengths. Recall that, for a realized spike history $X(T)$, the Fisher information functional can be written as
$\mathcal{J}[X]:=\int_{0}^{T}dt\sum_{k=1}^{N}\left[h_{k}^{\prime}(t,\theta)\frac{g_{k}^{\prime}(t)}{g_{k}(t)}\right]^{2}\rho_{k}(t),$ (7)
where $\rho_{k}(t)=g(u_{k}(t))$ and $g_{k}^{\prime}(t)=\left.\frac{dg}{du}\right|_{u=u_{k}(t)}$ (and similarly for higher derivatives). The Fisher information is $J=\langle\mathcal{J}[X]\rangle_{X(T)}$.
For any synaptic parameter $Z_{ij}\in\{w_{ij}^{0},\,U_{ij}\}$, differentiating the expectation $J(Z)=\langle\mathcal{J}[X]\rangle_{X(T)}$ yields the score-function identity
$\frac{\partial J}{\partial Z_{ij}}=\left\langle\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}+\mathcal{J}[X]\,\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{X(T)}.$ (8)
Because spikes are conditionally independent across neurons given the membrane potentials and the spike history, and because $Z_{ij}$ affects the membrane potential only through the postsynaptic neuron $i$, both terms in (8) can be expressed using an eligibility trace
$e_{ij}^{Z}(t):=\frac{\partial u_{i}(t)}{\partial Z_{ij}}.$ (9)
A direct differentiation of (7) gives
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\rho_{i}(t)\,\eta_{i}(t)\,e_{ij}^{Z}(t),$ (10)
where we introduced the information-weighting factor
$\eta_{i}(t):=\left[h_{i}^{\prime}(t,\theta)\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\right]^{2}\left(\frac{2g_{i}^{\prime\prime}(t)}{g_{i}^{\prime}(t)}-\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\right).$ (11)
For the exponential nonlinearity $g(u)=g_{c}e^{\beta(u-u_{c})}$, the ratio $g^{\prime}/g=\beta$ is constant and (11) simplifies to
$\eta_{i}(t)=\beta^{3}\,h_{i}^{\prime}(t,\theta)^{2}.$ (12)
Moreover, the score term admits the standard point-process form
$\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)=\int_{0}^{T}\left[dN_{i}(t)-\rho_{i}(t)\,dt\right]\,\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\,e_{ij}^{Z}(t),$ (13)
where $dN_{i}(t)$ is the increment of the counting process of neuron $i$ (equivalently, $\int_{0}^{T}\varphi(t)\,dN_{i}(t)=\sum_{f_{i}}\varphi(t_{i}^{f_{i}})$).
Equations (8)–(13) provide an exact representation of $\partial J/\partial Z_{ij}$ for finite recurrent coupling. A complete derivation is given in Appendix A.
3.1.2 Computability from network simulations
Importantly, the right-hand side of (8) is a function of quantities available along a single simulated trajectory $X(T)$: $\rho_{i}(t)$, $g_{i}^{\prime}(t)/g_{i}(t)$, and the eligibility trace $e_{ij}^{Z}(t)$. The trace $e_{ij}^{Z}(t)$ can be updated online during simulation by differentiating the membrane-potential dynamics with respect to $Z_{ij}$. For the exponential EPSP kernel $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the eligibility obeys a simple event-driven recursion (continuous decay between presynaptic spikes, with jumps at $t=t_{j}^{f}$ proportional to $\partial_{Z_{ij}}w_{ij}(t_{j}^{f})$), while the latter is obtained from the STP state variables and their parameter sensitivities. We provide explicit update equations for STD synapses in Appendix B. Thus, an unbiased Monte Carlo estimator of $\partial J/\partial Z_{ij}$ is obtained by averaging (8) over repeated simulations (or, in stationary/periodic regimes, by time-averaging after transients).
3.1.3 Weak-coupling reduction
While (8) is exact, it typically requires Monte Carlo averaging. Following [toyoizumi2006Fisher], we obtain a compact analytic form by assuming weak synaptic coupling, $w_{ij}^{0}\ll 1$ and consequently $w_{ij}(t)\ll 1$, and expanding around the baseline state $w_{ij}^{0}\equiv 0$. At $w=0$, the conditional intensity reduces to a deterministic function
$\nu_{i}^{0}(t):=g(h_{i}(t,\theta)),$ (14)
and $\mathcal{J}[X]$ becomes non-random; consequently, the score term in (8) vanishes at leading order. Using the compensation formula for inhomogeneous Poisson spiking at $w=0$ (Appendix C), the gradient reduces to
$\frac{\partial J}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\,\nu_{i}^{0}(t)\eta_{i}(t)\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\frac{\partial\langle w_{ij}(t^{\prime})\rangle_{X(t^{\prime})}}{\partial Z_{ij}}\,\nu_{j}^{0}(t^{\prime})\qquad(Z_{ij}=w_{ij}^{0},\,U_{ij}),$ (15)
where $\langle\cdot\rangle_{X(t)}$ denotes averaging over the baseline ($w=0$) spike statistics.
To characterize how short-term depression shapes the presynaptic contribution, we introduce normalized sensitivity functions
$f^{w_{0}}_{ij}(t)$ $:=\frac{1}{U_{ij}}\frac{\partial\langle w_{ij}(t)\rangle_{X(t)}}{\partial w_{ij}^{0}},$ (16a)
$f^{U}_{ij}(t)$ $:=\frac{1}{w^{0}_{ij}}\frac{\partial\langle w_{ij}(t)\rangle_{X(t)}}{\partial U_{ij}}.$ (16b)
Under weak coupling, these functions obey closed dynamics induced by the STD model (Appendix C.3):
$\dot{f}^{w_{0}}_{ij}$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+\nu_{j}^{0}(t)U_{ij}\right)f^{w_{0}}_{ij}(t),$ (17a)
$\dot{f}^{U}_{ij}$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+\nu_{j}^{0}(t)U_{ij}\right)f^{U}_{ij}(t)-\nu_{j}^{0}(t)U_{ij}f^{w_{0}}_{ij}(t).$ (17b)
Substituting (16) into (15) yields
$\frac{\partial J}{\partial w_{ij}^{0}}$ $=U_{ij}\int_{0}^{T}dt\,\nu_{i}^{0}(t)\eta_{i}(t)\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})f^{w_{0}}_{ij}(t^{\prime})\nu_{j}^{0}(t^{\prime}),$ (18a)
$\frac{\partial J}{\partial U_{ij}}$ $=w_{ij}^{0}\int_{0}^{T}dt\,\nu_{i}^{0}(t)\eta_{i}(t)\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})f^{U}_{ij}(t^{\prime})\nu_{j}^{0}(t^{\prime}).$ (18b)
3.1.4 “Hebbian” factorization and onset-sensitive presynaptic term
The weak-coupling gradient (18) admits a compact “Hebbian” form as a product of a postsynaptic factor and a filtered presynaptic factor:
$\frac{\partial J}{\partial Z_{ij}}$ $\propto\int_{0}^{T}dt\,\underbrace{C_{\mathrm{post},i}(t)}_{\text{post-component}}\underbrace{\left[\epsilon(t)*C_{\mathrm{pre},ij}^{Z}(t)\right]}_{\text{pre-component}},$ (19)
where $Z$ represents either $w_{0}$ or $U$. Here
$C_{\mathrm{post},i}(t):=\nu_{i}^{0}(t)\eta_{i}(t),$ (20)
and the effective presynaptic component is
$C_{\mathrm{pre},ij}^{Z}(t):=f^{Z}_{ij}(t)\nu_{j}^{0}(t).$ (21)
Thus, the learning rule detects the correlation between the post-synaptic information-bearing factor $C_{\mathrm{post},i}(t)$ and a filtered version of the effective presynaptic drive $C_{\mathrm{pre},ij}^{Z}(t)$. The crucial difference from standard Hebbian learning lies in the temporal dynamics of $C_{\mathrm{pre},ij}^{Z}(t)$: unlike conventional plasticity where the presynaptic term is simply the firing rate $\nu_{j}^{0}(t)$, here the rate is modulated by the sensitivity function $f^{Z}_{ij}(t)$, which depends on the recent history of presynaptic activity through STD.
3.1.5 Response to Step and Sinusoidal Inputs.
To gain intuition, we first examine the behavior of these terms under constant stimulation (Figure 1 A-C). Both $f^{w_{0}}_{ij}(t)$ and $f^{U}_{ij}(t)$ decay exponentially upon the onset of presynaptic firing. However, $f^{U}_{ij}(t)$ exhibits a stronger suppression and a more prolonged transient response compared to $f^{w_{0}}_{ij}(t)$. This suggests that the learning of the release probability $U$ is more sensitive to the onset of activity bursts than the learning of the baseline weight $w_{0}$.
This onset-sensitivity becomes even more apparent under sinusoidal stimulation (Figure 1 D-F). Since $f^{Z}_{ij}(t)$ decreases as the firing rate $\nu_{j}^{0}(t)$ increases (an anti-phase relationship), the effective presynaptic term $C^{Z}_{ij}(t)=f^{Z}_{ij}(t)\nu_{j}^{0}(t)$ does not peak when the firing rate is maximal. Instead, it peaks during the rising phase (onset) of the firing rate, before significant depression has accumulated. Consistent with the step response, this effect is more pronounced for the release probability parameter: $C^{U}_{ij}(t)$ peaks earlier and is more strongly suppressed than $C^{w_{0}}_{ij}(t)$.
Figure 1: Response of STP-derived sensitivity functions to step / sinusoidal inputs. A-C. Response to step input. A. Presynaptic firing rate $\nu_{j}(t)$. B. Dynamics of sensitivity functions $f^{Z}$. Note that $f^{U}$ (orange) decays more strongly than $f^{w_{0}}$ (blue). C. Dynamics of the effective presynaptic contribution $C(t)=f(t)\nu_{j}(t)$. D-F. Response to sinusoidal input. The effective presynaptic term (F) peaks during the rising phase of the input (D), illustrating the onset-detection property of the learning rule. Parameters: $\tau_{d}=0.5$ s, $U=0.15$.
We further quantified this temporal shift by analyzing the linear response to weak sinusoidal modulation. Figure 2 illustrates the frequency response of the effective presynaptic term $C^{Z}(t)$, summarized by its complex gain $H_{C_{\mathrm{pre}}}^{Z}(\omega)$, whose magnitude and phase give the amplitude gain and phase shift relative to the input (see Appendix D.2 for details).
Figure 2: Frequency-domain response of the effective presynaptic term $C^{Z}(t)$. Response to a modulated rate $\nu(t)=\nu_{0}+\delta\nu\cos(\omega t)$. Upper. Amplitude gain $|H_{C_{pre}}^{Z}(\omega)|$. Lower. Phase lag $\phi_{C_{pre}}^{Z}(\omega)=\arg H_{C_{pre}}^{Z}(\omega)$. Note the positive phase shift (lead) in the intermediate frequency range, indicating sensitivity to the input onset. The lead is more pronounced for $U$ (orange) than for $w_{0}$ (blue). Lines show the predictions by linear response theory, and dots show the results by numerical simulations. Parameters: $\nu_{0}=10$ Hz, $\tau_{d}=0.5$ s, $U=0.15$.
Three key regimes are observed:
  1. 1.
    Low Frequency: The system behaves quasi-statically. There is no phase lag relative to the input, and the amplitude matches the steady-state derivative.
  2. 2.
    High Frequency: The dynamics are low-pass filtered, dampening the oscillatory response.
  3. 3.
    Intermediate Frequency: A distinct resonance emerges. The phase lead of $C^{Z}(t)$ relative to the input $\nu(t)$ becomes positive, confirming that the learning rule selectively highlights the rising phase of the input.
Crucially, the phase lead for $U$ is consistently larger than for $w_{0}$ across relevant frequencies. Also, the frequencies at which these responses peak ($\omega_{*}$) differ between parameters, implying that $w_{0}$ and $U$ are tuned to extract different temporal features of the input statistics. A more nuanced characterization of how these gains and phase relationships depend on frequency, baseline rate, and STP parameters is presented in Appendix D.2.
3.1.6 Implications for Learning.
A natural implication of our Fisher-information optimization analysis in the above sections is the emergence of anti-causal feedback connectivity patterns that could support reverse replay–like activity. At the level of individual synapses, the postsynaptic factor $\nu_{i}^{0}(t)\eta_{i}(t)$ represents the local Fisher-information contribution and typically scales with the postsynaptic firing rate. Consequently, the learning rule reinforces connections where the onset of presynaptic activity (high $C^{Z}_{ij}$) predicts high postsynaptic information (high $\nu_{i}\eta_{i}(t)$). Because the presynaptic term is effectively phase-advanced (peaking before the rate maximum), this learning rule naturally favors anti-causal associations, i.e., connections from neurons that fire later to neurons that fired earlier in the stimulus-evoked sequence. This tendency is consistent for both the release probability $U_{ij}$ and the baseline weight $w^{0}_{ij}$, but the stronger phase lead in $U_{ij}$ suggests that the optimization of synaptic reliability is particularly driven by rapid temporal changes in the inputs (approximately the temporal derivative), potentially enabling the network to encode dynamic stimulus features more efficiently than static firing rates. At the network level, such anti-causal feedback projections imply that neurons with later preferred positions or phases tend to drive neurons with earlier preferences. This backward connectivity motif has been shown to support reverse replay in hippocampal place-cell models [haga2018Recurrent].
3.2 Case Study: Traveling Waves on a Circle
To better understand the circuit-level consequences of the learning rules, we consider neurons arranged on a circle receiving traveling-wave input. In this case study, we first use the weak-coupling approximation to obtain analytic intuition for the phase structure of the Fisher-information gradients under rotational symmetry. We then quantify learning-induced changes in neural representations and Fisher information in the full recurrent network using the exact (simulation-based) gradient estimator derived in Sec. 3.1.1.
3.2.1 Settings
We consider neurons uniformly distributed on a circle, taking the limit $N\rightarrow\infty$. Each neuron is specified by its angular position $z$ on the circle, which physiologically corresponds to its preferred orientation. Neurons receive traveling wave input $h(z,t)=h(\omega t-z)$ with constant velocity and shape. We assume that the learned parameters possess rotational symmetry, such that $w_{0}(z,z^{\prime})=w_{0}(\zeta)$ and $U(z,z^{\prime})=U(\zeta)$, where $\zeta=z-z^{\prime}$ represents the phase difference between presynaptic position $z^{\prime}$ and postsynaptic position $z$.
In the following analysis, we specifically consider the input $h(z,t)=A[\cos(\omega t-z)-\cos\theta_{c}]_{+}$, which represents:
  • Spatially: a bump of width $2\theta_{c}$ centered at $z=\omega t$, rotating with angular velocity $\omega$
  • Temporally: a bump of duration $2\theta_{c}/\omega$ centered at $t=(z+2\pi n)/\omega$, with frequency $\omega/2\pi$
The encoded parameter is $\theta_{c}$, which determines the bump width.
For finite recurrent coupling, the microscopic gradients $\partial J/\partial Z(z,z^{\prime})$ (with $Z\in\{w_{0},U\}$) can be estimated from network simulations via the exact score-function/eligibility-trace identity (Eq. (8)). Rotational symmetry then implies that the functional derivative with respect to the profile $Z(\zeta)$ is obtained by averaging over all pre–post pairs that share the same phase offset:
$\frac{\delta J}{\delta Z(\zeta)}=\frac{1}{2\pi}\int_{-\pi}^{\pi}dz\;\left.\frac{\partial J}{\partial Z(z,z^{\prime})}\right|_{z^{\prime}=z-\zeta}\qquad(Z=w_{0},U).$ (22)
We use this exact estimator when assessing the learning-induced changes in network representations and information transmission.
To obtain a transparent expression for the phase dependence of the gradients, we also consider the weak-coupling approximation developed above. In this limit, the profile gradients reduce to:
$\begin{split}\frac{\delta J}{\delta w_{0}(\zeta)}&=U(\zeta)\int_{-\pi}^{\pi}\frac{1}{2\pi}dz\int_{0}^{T}dt\;\nu^{0}(z,t)\eta^{0}(z,t)\\ &\quad\times\int_{0}^{t}dt^{\prime}\;\epsilon(t-t^{\prime})\,f^{w_{0}}(t^{\prime};z,z-\zeta)\,\nu^{0}(z-\zeta,t^{\prime}),\end{split}$ (23a)
$\begin{split}\frac{\delta J}{\delta U(\zeta)}&=w_{0}(\zeta)\int_{-\pi}^{\pi}\frac{1}{2\pi}dz\int_{0}^{T}dt\;\nu^{0}(z,t)\eta^{0}(z,t)\\ &\quad\times\int_{0}^{t}dt^{\prime}\;\epsilon(t-t^{\prime})\,f^{U}(t^{\prime};z,z-\zeta)\,\nu^{0}(z-\zeta,t^{\prime}).\end{split}$ (23b)
Here $\nu^{0}(z,t)=g(h(z,t))$, and $\eta^{0}(z,t)$ is the corresponding information-weighting factor (Eq. (11) evaluated at $u=h$; for the exponential nonlinearity, Eq. (12)). The auxiliary functions $f^{w_{0}}(t;z,z^{\prime})$ and $f^{U}(t;z,z^{\prime})$ are the weak-coupling sensitivity functions (Eq. (16)) and obey the same dynamics as Eq. (17) with the substitutions $\nu_{j}^{0}(t)\to\nu^{0}(z^{\prime},t)$ and $U_{ij}\to U(z-z^{\prime})$.
3.2.2 Optimization of synaptic parameters
In this subsection, we use the weak-coupling gradients in Eq. (23) to obtain an analytic prediction for how the optimal profiles $U(\zeta)$ and $w_{0}(\zeta)$ depend on the phase offset $\zeta$.
To visualize how the “Hebbian” decomposition (Sec. 3.1.4) manifests in the present traveling-wave setting, Figure 3 plots the local postsynaptic factor $C_{\mathrm{post}}(z,t)=\nu^{0}(z,t)\eta^{0}(z,t)$ and the effective presynaptic factor $C_{\mathrm{pre}}^{Z}(t;z,\zeta)=f^{Z}(t;z,z-\zeta)\nu^{0}(z-\zeta,t)$ (for $Z\in\{w_{0},U\}$), for a representative anti-causal offset ($\zeta>0$) where postsynaptic activity precedes presynaptic activation and the temporal overlap between pre- and postsynaptic factors is large. Consistent with the preceding analysis, short-term depression makes $C_{\mathrm{pre}}^{Z}$ dominated by stimulus onset. Notably, the $U$ factor $C_{pre}^{U}(t)$ can become negative during sustained stimulation (Fig. 3): raising $U$ may paradoxically suppress synaptic transmission by accelerating resource depletion, highlighting the distinct sensitivity of release-probability modulation. The gradient contribution is therefore controlled primarily by how strongly the onset-weighted presynaptic drive overlaps with $C_{\mathrm{post}}$.
Figure 3: Analysis of pre- and postsynaptic factors (for representative anti-causal pair). Pre- and postsynaptic neurons receive the same traveling-wave external input $h(z,t)=A[\cos(\omega t-z)-\cos\theta_{c}]_{+}$ at different phases. The plots illustrate a representative anti-causal offset ($\zeta>0$; postsynaptic activation precedes presynaptic activation), where the pre/post overlap is large. The STP factors $f^{w_{0}}$ and $f^{U}$ decay rapidly with increasing firing rate, so the presynaptic contributions $C_{\mathrm{pre}}^{w_{0}}$ and $C_{\mathrm{pre}}^{U}$ are dominated by stimulus onset. The contribution to the Fisher-information gradient is determined by the temporal overlap between $C_{\mathrm{pre}}$ and the postsynaptic term $C_{post}$. Parameters: $\tau_{d}=0.5~\mathrm{s}$, $\beta=2.0$, $g_{M}=10.0$, $u_{c}=1.0$, $\tau_{s}=0.01~\mathrm{s}$, $A=3.0$, $\theta_{c}=0$.
Under the same weak-coupling approximation, we can display the optimal profiles explicitly in close analogy to [toyoizumi2006Fisher]. For $w_{0}(\zeta)$, we impose a zero-mean (balance) constraint and a fixed variance (synaptic cost), which yields an optimal profile that is proportional to the gradient shape:
$w_{0}(\zeta)\propto\frac{\delta J}{\delta w_{0}(\zeta)},$ (24)
up to an additive constant and overall scaling set by the constraints. For $U(\zeta)$, we impose a fixed mean release-probability budget $\frac{1}{2\pi}\int_{-\pi}^{\pi}d\zeta\,U(\zeta)=\bar{U}$ (and the bound $U(\zeta)\in[0,1]$), so that the optimum is selected by the level sets of $\delta J/\delta U(\zeta)$ under this budget.
Figure 4 visualizes $\delta J/\delta U(\zeta)$ as a function of $\zeta$ and $U$; the white contours indicate optimal solutions under the constant-sum constraint. The resulting optimal $U(\zeta)$ (Figure 5A) is strongly temporally asymmetric: release probability is maximal when the presynaptic phase lags the postsynaptic phase ($\zeta>0$), and remains low when presynaptic activity leads ($\zeta<0$). Thus, the weak-coupling optimum predicts an anti-causal bias in the learned connectivity. Optimizing $w_{0}(\zeta)$ under this optimized $U(\zeta)$ produces a similarly asymmetric $w_{0}(\zeta)$ (Figure 5B,C), whereas holding $U(\zeta)$ fixed substantially attenuates the asymmetry (Figure 5B,C), highlighting that plasticity of release probability is essential for expressing the full temporal bias.
Figure 4: Gradient of Fisher information with respect to $U(\zeta)$. Heat map of $\frac{\delta J}{\delta U(\zeta)}$ as a function of phase difference $\zeta$ and $U$. White contour lines indicate optimal solutions under the constant-sum constraint. The red contour marks the boundary where $\frac{\delta J}{\delta U(\zeta)}=0$. Parameters are identical to Figure 3.
Figure 5: Optimal synaptic profiles. A. Optimal release probability $U(\zeta)$. B. Optimal baseline weight $w_{0}(\zeta)$. Green, solid: optimization with plastic $U(\zeta)$. Yellow, dashed: optimization with non-plastic $U$. C. Optimal effective weight $w(\zeta)=w_{0}(\zeta)U(\zeta)$. Purple, solid: optimization with plastic $U(\zeta)$. Red, dashed: optimization with non-plastic $U$. Optimal $U(\zeta)$ is derived under the constraint $\frac{1}{2\pi}\int_{-\pi}^{\pi}d\zeta\,U(\zeta)=0.15$. In the non-plastic condition, $U$ is fixed to 0.15. Other parameters are identical to Figure 4.
3.2.3 Constraint-dependent modulation of temporal asymmetry
We next examine how the temporal asymmetry of the optimal weights depends on constraints on release probability and on the strength of external inputs (Figure 6; see also Supplementary Figures S7, S8). Allowing higher release probabilities or stronger inputs consistently enhances the anti-causal bias of both $U(\zeta)$ and $w(\zeta)$, whereas low release probabilities lead to nearly symmetric potentiation and depression profiles.
These findings indicate that the effective learning rule is not fixed, but instead is modulated by global constraints on synaptic reliability and input drive. In particular, regimes with low release probability naturally suppress anti-causal biases and recover almost symmetric weight updates. In this low-$U$ limit, short-term dynamics of synapses are effectively muted, so that resulting updates closely resemble the case without STP. Whrease regimes with high release probability amplify anti-causal associations. This constraint-dependent modulation provides a plausible mechanism for state-dependent changes in synaptic learning rules, which we further relate to sleep–wake transitions in the Discussion.
Figure 6: Low release probabilities attenuate temporal asymmetry. Comparison of optimal $w(\zeta)$ for different constraints on the mean release probability. Solid: mean $U(\zeta)=0.15$. Dashed: mean $U(\zeta)=0.03$. When release probability is constrained to low values, temporal asymmetry in the learning rule is substantially reduced, approaching a nearly symmetric profile. Other parameters are identical to Figure 4.
3.2.4 Changes in Neural Representation
To assess how the learning rules reshape neural representations, we examine the recurrent synaptic input and the resulting firing-rate profile (Figure 7). After learning, the synaptic input current is concentrated in periods of postsynaptic activity and remains close to zero when the postsynaptic neuron is inactive. Within the active window, the input becomes biased toward the latter half of the bump. Correspondingly, the firing-rate bump shifts its peak to later times and exhibits asymmetry between onset and offset. These changes indicate that learning redistributes synaptic gain within the bump, selectively emphasizing later phases of the response instead of uniformly amplifying activity.
We next compare different optimization conditions. When both $U$ and $w_{0}$ are plastic, the representation exhibits the clearest temporal asymmetry and the largest contrast between active and inactive periods. If $U$ is held fixed and only $w_{0}$ is optimized, the synaptic input still grows during bump activity, but the bias toward the latter half and the resulting temporal asymmetry are substantially weakened. Conversely, when $U$ remains plastic but $w_{0}$ is fixed, the bump becomes more asymmetric in time, but the overall modulation amplitude is reduced. Finally, in networks without STP dynamics, synaptic input lacks temporal asymmetry and instead develops negative components during quiescent periods. This inhibitory input outside the postsynaptic activity window, however, barely affects the bump profile, indicating that recurrent resources are spent without effective modulation of the neural activity. In contrast, STP-based models confine synaptic input to task-relevant epochs, thereby achieving more efficient representations for a given synaptic cost.
Figure 7: Changes in neural representation induced by synaptic optimization. Upper panel: recurrent synaptic input $\int_{-\pi}^{\pi}d\zeta\int dt^{\prime}\,\epsilon(t-t^{\prime})w(\zeta,t)\,\nu^{0}(-\zeta)$ to a representative neuron. Lower panel: corresponding firing-rate bump $\nu(t)$. Blue: joint optimization of $U$ and $w_{0}$ (plastic STP and baseline weight). Purple: optimization of $w_{0}$ with fixed $U$. Light blue: optimization of $U$ with fixed $w_{0}$. Green: optimization of a static synapse $w$ without STP dynamics. Black dash (in lower panel): original firing rate $\nu_{0}(t)$.
3.2.5 Fisher Information Enhancement
We next quantify how these representation changes affect the Fisher information. For small synaptic modifications, the change in Fisher information can be approximated as
$\Delta J\approx\int_{0}^{T}dt\int_{-\pi}^{\pi}dz\,h^{\prime}(z,t)\left(\frac{g^{\prime 0}(z,t)}{g^{0}(z,t)}\right)^{2}\Delta\nu(z,t),$ (25)
where $\Delta\nu(z,t)$ denotes the learning-induced change in firing rate.
Because an overall increase in firing rate trivially elevates Fisher information in the linear-response regime, we compare different optimization schemes under a fixed magnitude of membrane-potential perturbation. Specifically, we constrain the total deviation
$\mathcal{V}=\int_{0}^{T}dt\int_{-\pi}^{\pi}dz\,[\Delta u(z,t)]^{2},$ (26)
where $\Delta u(z,t):=u(z,t)-h(z,t)$ is the change in membrane potential induced by recurrent input. Figure 8 shows the resulting Fisher-information gain $\Delta J/J^{0}$ for each condition, normalized by $\mathcal{V}$.
Among synapses with and without STP, Fisher information is maximally enhanced when both $U$ and $w_{0}$ are plastic. In this case, associative learning jointly shapes presynaptic release dynamics and postsynaptic gain, confining synaptic input to the bump while suppressing it during quiescent periods. This selective gating allows the network to approach an optimal trade-off between information enhancement and deviation from the original activity pattern, effectively maximizing information gain per unit of metabolic or homeostatic cost.
Figure 8: Fisher information enhancement under different optimization schemes. Relative change in Fisher information $\Delta J/J^{0}$ across optimization conditions, normalized by the total membrane-potential fluctuation $\mathcal{V}$. Synapses with STP achieve the largest information gain when both $U$ and $w_{0}$ are plastic.
4 Discussion
We have demonstrated that extending Fisher information optimization to synapses with Tsodyks–Markram dynamics yields learning rules that preferentially strengthen anti-causal connections, where presynaptic neurons lag behind postsynaptic neurons. This temporal asymmetry emerges from the interaction between associative learning and short-term depression, creating neural representations that emphasize stimulus offsets through ramp-up dynamics. The degree of temporal asymmetry depends critically on constraints governing vesicle release probability. As we discuss below, this combination of anti-causal connectivity and release-probability constraints provides a plausible computational account of state-dependent differences in hippocampal replay directionality observed during wakefulness versus sleep.
4.1 Pre- versus postsynaptic plasticity: Computational division of labor
The segregation of synaptic plasticity mechanisms between pre- and postsynaptic sites has long served as a guiding principle for understanding synaptic computation. Classical long-term potentiation is mediated primarily by postsynaptic mechanisms, notably changes in AMPA receptor trafficking and expression [malenka2004LTP]. By contrast, short-term plasticity has typically been modeled as a predominantly presynaptic phenomenon governed by vesicle dynamics and calcium-dependent modulation of release probability [zucker2002Shortterm, regehr2012ShortTerm]. This division of labor has been interpreted as separating associative learning (postsynaptic, detecting pre-post correlations) from adaptive filtering (presynaptic, responding only to presynaptic history).
Recent theoretical work has argued that this pre–post division reflects a deeper computational separation rather than a purely biophysical contingency. Postsynaptic mechanisms, operating on longer timescales, are well suited for stable memory storage and the formation of neural assemblies that encode persistent representations [fusi2005Cascade, zenke2015Diverse]. Presynaptic mechanisms, with their rapid dynamics, can implement adaptive filters that extract temporal features from spike trains, enhance signal-to-noise ratios, and maintain homeostatic balance in network activity [abbott2004Synaptic, fortune2001Shortterm].
The discovery of associative short-term plasticity by Ucar et al. [ucar2021Mechanical] challenges this traditional dichotomy. Their experiments demonstrate that presynaptic release probability can be modulated via postsynaptic spine enlargement, in a manner that requires coincident pre- and postsynaptic activity. This mechanism therefore combines the associative detection capabilities conventionally attributed to postsynaptic plasticity with the rapid, history-dependent filtering characteristic of presynaptic dynamics. Our theoretical framework indicates that such a hybrid rule is not simply “LTP made faster”, but gives rise to distinct forms of temporal selectivity.
In our model, the interaction between associative learning rules and short-term depression generates temporal selectivity that is difficult to realize with static synaptic weights alone. When release probability $U$ is modulated associatively, the resulting synaptic dynamics become sensitive to the temporal structure of presynaptic spike trains in a context-dependent manner. Increased $U$ amplifies responses at the onset of presynaptic bursts, while depletion suppresses sustained firing, effectively implementing a high-pass filter whose characteristics are shaped by the postsynaptic activation patterns. This provides a mechanism for rapid, selective enhancement of behaviorally relevant temporal features, as signaled by postsynaptic activation.
4.2 Reverse replay and state-dependent plasticity
[…]
This mechanism aligns with and extends the model proposed by [haga2018Recurrent], which demonstrated that symmetric STDP combined with short-term depression can generate backward connectivity supporting reverse replay. In our framework, such connectivity patterns arise from an information-theoretic optimization principle: a backward bias is not an incidental artifact of a particular plasticity rule, but an expected outcome when information transmission is optimized under appropriate constraints.
A key result is that the strength of temporal asymmetry depends on the total release-probability budget $\sum_{j}U_{ij}$. Under tight constraints (small $\sum_{j}U_{ij}$), the optimal learning rule is nearly symmetric in time, with balanced potentiation and depression. As the constraint is relaxed (larger $\sum_{j}U_{ij}$), the rule becomes increasingly asymmetric, with dominant anti-causal potentiation. In this sense, the extent of reverse connectivity is controlled by how liberally presynaptic resources can be allocated to support plastic changes.
We interpret this “release-probability budget” as an effective parameter that can be influenced by physiological mechanisms such as neuromodulation. Neuromodulators vary strongly across behavioral states and sleep stages and modulate presynaptic properties, including short-term dynamics [hasselmo2006Role, sara2009Locus]. Recent work further underscores that presynaptic resources are regulated at the behavioral-state level: [wu2025Presynaptic] showed across multiple circuits in Drosophila and mice that presynaptic release probability decreases with prolonged wakefulness and recovers after sleep, and that bidirectional manipulation of release probability bidirectionally alters sleep need. Complementary studies in Drosophila demonstrate that presynaptic active zone composition, which closely tracks release probability, is remodeled as a function of sleep pressure [huang2020Presynaptic]. Although these findings do not address hippocampal replay directly, they establish that presynaptic resource constraints can indeed be tuned in a state-dependent manner, consistent with the key premise of our model.
Within this interpretation, our framework offers a hypothesis for why replay directionality depends on behavioral state. Hippocampal circuits are expected to possess a baseline forward bias in connectivity, for example through classical causal STDP rules during spatial exploration, which favors forward replay by default. Our analysis shows that associative short-term plasticity optimized under a generous presynaptic budget superimposes an additional reverse component onto this baseline. When presynaptic resources are effectively abundant, the anti-causal contribution can partially compensate or even outweigh the baseline forward asymmetry, yielding robust reverse or bidirectional replay. When the budget is more stringent, the reverse component weakens and the baseline forward structure dominates, resulting in predominantly forward replay.
This picture dovetails with empirical observations that reverse replay is especially prominent during quiet wakefulness following spatial experience and reward delivery, whereas forward replay is more prevalent during sleep sharp-wave ripples [foster2006Reverse, diba2007Forward, ambrose2016Reverse, buzsaki2015Hippocampal, olafsdottir2018Role]. In our framework, these regimes correspond to different effective presynaptic budgets acting on top of a fixed forward backbone: more generous budgets promote the emergence of reverse replay that may facilitate rapid behavioral adaptation and reward credit assignment [foster2006Reverse], whereas more stringent budgets yield connectivity that expresses predominantly forward replay and may better support stabilization and generalization of memories during sleep [gais2004Declarative, rasch2013Sleeps, buzsaki2015Hippocampal, olafsdottir2018Role]. This proposal links state-dependent neuromodulation of presynaptic function to a quantitative, information-theoretic control knob for the balance between forward and reverse replay, and thereby frames reverse sequences not as an oddity of hippocampal dynamics, but as a predicted feature of efficient temporal coding under varying resource constraints.
4.3 Frequency-Dependent Phase Selectivity in Pre- and Postsynaptic Channels
Throughout this chapter, we have emphasized that short-term plasticity (STP) acts as a detector of presynaptic onset. While useful for intuition, this view is an approximation: the strength of onset-driven plasticity depends on input frequency and amplitude, baseline firing rates, release probability, membrane potential, and STP time constants.
Our linear-response analysis of $C^{Z}(t)$ (Figure 2; Appendix D.2) confirms that onset selectivity is robust, but the phase of maximal plasticity shifts with input frequency and the neuron–synapse operating point. Thus, STP implements a frequency-dependent phase preference rather than a simple onset detector. Furthermore, if STP parameters such as release probability $U$ are modified by learning or neuromodulation, phase selectivity shifts accordingly, generating heterogeneity across synapses in which temporal relationships are preferentially learned.
[…]
The combination of presynaptic dynamics (frequency- and phase-dependent filtering via STP) and postsynaptic phase sensitivity (encoding information content) thus provides a two-factor mechanism for shaping synaptic structure.
Within the linear-response regime, this framework also extends to inputs with multiple frequency components. One can then characterize how STP and phase-dependent postsynaptic sensitivity jointly sculpt plasticity across a spectrum of frequencies and phases, treating STP as part of a general temporal signal-processing mechanism.
4.4 Assumptions on Short-Term Facilitation
In subsection 2.2, we neglected short-term facilitation (STF) and modeled short-term plasticity (STP) solely as short-term depression (STD). Experimentally, the facilitation time constant $\tau_{f}$ is typically on the order of tens to a few hundred milliseconds, whereas the recovery time constant from depression $\tau_{d}$ is on the order of several hundred milliseconds up to seconds[zucker2002Shortterm, regehr2012ShortTerm]. Because the associative STP we aim to model is induced on a timescale of at least seconds [kasai2023Mechanical, ucar2021Mechanical], the interaction we primarily need to capture is that between associative plasticity and the slowly recovering depressive component.
Furthermore, in associative forms of STP the release probability $U$ after induction is often driven close to one, which further reduces the effective impact of STF. Nevertheless, before induction, when the baseline release probability is low, STF may transiently influence synaptic responses. A more refined analysis could therefore extend our framework to include STF explicitly and quantify how it shapes the early dynamics of associative STP.
4.5 Limitations and future directions
Several limitations of our current framework warrant discussion and motivate future extensions. First, our analysis employs a perturbative approach, examining small deviations from baseline activity patterns. This limitation prevents us from capturing the full recursive dynamics where recurrent inputs reshape representations, which in turn modify learning rules, creating complex feedback loops. The interaction between synaptic dynamics and neural dynamics likely becomes particularly important when depression timescales approach those of neural integration, potentially leading to rich computational regimes not captured by our linear analysis [hansel2013Shortterm, ocker2015SelfOrganization].
[…]
Third, we model spiking as an inhomogeneous Poisson process without refractoriness or reset mechanism, which yields a tractable link between filtered synaptic input and instantaneous firing rate. This approximation is most appropriate in regimes where spiking is highly irregular and the relevant stimulus/synaptic timescales are slow compared with the refractory timescale, such that spike-history effects contribute only weak corrections to rate-based statistics and information measures. Outside this regime, however, real neurons exhibit absolute/relative refractoriness and other spike-history dependencies that induce negative serial correlations and can interact with short-term depression in complex ways [schwalger2017Theory]. These history-dependent effects may therefore modify both Fisher information and the optimal plasticity updates predicted by our framework. A natural future direction is to extend the model to a more general conditional-intensity (escape-rate/GLM) formulation with an explicit spike-history term (or renewal/dead-time dynamics), and re-derive the Fisher information gradients under this enriched point-process description.
[…]
Despite these limitations, our framework is readily extensible to address many of these challenges. The mathematical formalism developed here can accommodate different models of presynaptic dynamics, including short-term facilitation-dominant regimes, which may yield qualitatively different learning rules and representation strategies. Similarly, extending the analysis to alternative objective functions – such as mutual information [toyoizumi2005Generalized, toyoizumi2005Spiketiming, pfister2006Optimal], sparse coding [olshausen1996Emergence], or predictive coding [rao1999Predictive] – would be promising within our framework.
5 Conclusion
We have shown that optimizing Fisher information in synapses with short-term dynamics yields learning rules that preferentially strengthen anti-causal connections, where presynaptic neurons lag behind postsynaptic targets. This temporal asymmetry arises from the interaction between associative learning and synaptic depression, creating representations that emphasize stimulus offsets and naturally support reverse replay. The strength of this asymmetry depends on release probability constraints, potentially explaining state-dependent differences in replay directionality between wakefulness and sleep. These results suggest that associative short-term plasticity is not merely“faster LTP” but enables distinct computational capabilities through the combination of correlation detection and dynamic filtering that static weight changes alone cannot achieve.
Appendix A Exact gradient formula
This appendix derives the exact gradient expression (8) and the explicit forms (10)–(13) used in Section 3.1.1.
A.1 Point-process likelihood and the score-function identity
Let $N_{i}(t)$ denote the counting process of neuron $i$ and $dN_{i}(t)$ its increment. Under our conditional independence assumption, the joint likelihood of the population spike history $X(T)$ factorizes across neurons given the conditional intensities $\rho_{i}(t)=g(u_{i}(t))$:
$\log P(X(T)\mid Z)=\sum_{k=1}^{N}\left[\int_{0}^{T}\log\rho_{k}(t)\,dN_{k}(t)-\int_{0}^{T}\rho_{k}(t)\,dt\right].$ (A.1)
(Equivalently, $\int_{0}^{T}\varphi(t)\,dN_{k}(t)=\sum_{f_{k}}\varphi(t_{k}^{f_{k}})$ for any test function $\varphi$.)
For any scalar objective written as an expectation over spike histories,
$J(Z)=\left\langle\mathcal{J}[X,Z]\right\rangle_{X(T)}=\int\mathcal{J}[X,Z]\,P(X(T)\mid Z)\,dX,$ (A.2)
differentiation under the integral gives the standard score-function (likelihood-ratio) identity
$\frac{\partial J}{\partial Z}=\left\langle\frac{\partial\mathcal{J}[X,Z]}{\partial Z}+\mathcal{J}[X,Z]\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}.$ (A.3)
Applying (A.3) to $Z=Z_{ij}$ yields (8).
Differentiating (A.1) gives
$\frac{\partial}{\partial Z}\log P(X(T)\mid Z)=\sum_{k=1}^{N}\int_{0}^{T}\left[dN_{k}(t)-\rho_{k}(t)\,dt\right]\,\frac{\partial}{\partial Z}\log\rho_{k}(t).$ (A.4)
In our model $\rho_{k}(t)=g(u_{k}(t))$, hence
$\frac{\partial}{\partial Z}\log\rho_{k}(t)=\frac{g_{k}^{\prime}(t)}{g_{k}(t)}\,\frac{\partial u_{k}(t)}{\partial Z},\qquad g_{k}^{\prime}(t):=\left.\frac{dg}{du}\right|_{u=u_{k}(t)}.$ (A.5)
For the synapse-specific parameter $Z_{ij}\in\{w_{ij}^{0},U_{ij}\}$, the membrane potential $u_{k}(t)$ depends on $Z_{ij}$ directly only when $k=i$. Therefore (A.4)–(A.5) reduce to
$\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)=\int_{0}^{T}\left[dN_{i}(t)-\rho_{i}(t)\,dt\right]\,\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)},$ (A.6)
which is (13) in the main text.
A.2 Pathwise derivative of the Fisher-information functional
We next compute $\partial\mathcal{J}[X]/\partial Z_{ij}$ for the Fisher-information functional (7). Define
$A_{k}(t):=h_{k}^{\prime}(t,\theta)\frac{g_{k}^{\prime}(t)}{g_{k}(t)},\qquad\rho_{k}(t)=g_{k}(t):=g(u_{k}(t)).$ (A.7)
Then (7) is
$\mathcal{J}[X]=\int_{0}^{T}dt\sum_{k=1}^{N}A_{k}(t)^{2}\,\rho_{k}(t).$ (A.8)
For a fixed spike history $X(T)$, the dependence on $Z_{ij}$ enters only through the membrane potentials $u_{k}(t)$, and hence through $g_{k}(t),g_{k}^{\prime}(t),g_{k}^{\prime\prime}(t)$. Differentiating with the chain rule gives
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\frac{\partial}{\partial u_{i}}\left(h_{i}^{\prime}(t,\theta)^{2}\,\frac{g_{i}^{\prime}(t)^{2}}{g_{i}(t)}\right)\,\frac{\partial u_{i}(t)}{\partial Z_{ij}}.$ (A.9)
A short calculation yields
$\frac{\partial}{\partial u}\left(\frac{g^{\prime}(u)^{2}}{g(u)}\right)=\frac{g^{\prime}(u)^{2}}{g(u)}\left(\frac{2g^{\prime\prime}(u)}{g^{\prime}(u)}-\frac{g^{\prime}(u)}{g(u)}\right).$ (A.10)
Using $\rho_{i}(t)=g_{i}(t)$ and defining $\eta_{i}(t)$ by (11), we obtain
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\rho_{i}(t)\,\eta_{i}(t)\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)}.$ (A.11)
This establishes (10).
A.3 A note on unbiased estimation and variance reduction
Combining (8), (A.6), and the pathwise term yields an unbiased estimator from simulated trajectories:
$\begin{split}\widehat{\frac{\partial J}{\partial Z_{ij}}}&=\frac{1}{M}\sum_{m=1}^{M}\Bigg[\int_{0}^{T}dt\;\rho_{i}^{(m)}(t)\eta_{i}^{(m)}(t)e_{ij}^{Z,(m)}(t)\\ &+\mathcal{J}[X^{(m)}]\int_{0}^{T}\left(dN_{i}^{(m)}(t)-\rho_{i}^{(m)}(t)\,dt\right)\frac{g_{i}^{\prime,(m)}(t)}{g_{i}^{(m)}(t)}e_{ij}^{Z,(m)}(t)\Bigg].\end{split}$ (A.12)
As is common for score-function estimators, the second term can have high variance. A standard control variate is to replace $\mathcal{J}[X^{(m)}]$ by $\mathcal{J}[X^{(m)}]-b$ with a constant $b$ (e.g., an online running mean), which preserves unbiasedness because $\langle\partial_{Z}\log P\rangle=0$:
$\left\langle b\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}=b\,\frac{\partial}{\partial Z}\int P(X(T)\mid Z)\,dX=0.$ (A.13)
Appendix B Eligibility traces for STD synapses
This appendix provides explicit online update equations for the eligibility trace $e_{ij}^{Z}(t)=\partial u_{i}(t)/\partial Z_{ij}$ for $Z_{ij}\in\{w_{ij}^{0},U_{ij}\}$, under the Tsodyks–Markram STD-only dynamics used in the main text.
B.1 Eligibility as a filtered, parameter-weighted presynaptic spike train
Write the presynaptic spike train as $x_{j}(t)=\sum_{f}\delta(t-t_{j}^{f})$ and define the weighted spike train
$y_{ij}(t):=\sum_{f}w_{ij}(t_{j}^{f})\,\delta(t-t_{j}^{f})=w_{ij}(t^{-})\,x_{j}(t),$ (B.1)
where $t^{-}$ indicates evaluation immediately before a spike at time $t$. Then the synaptic contribution from neuron $j$ to $i$ can be written as a convolution
$u_{i}(t)=h_{i}(t,\theta)+\sum_{k=1}^{N}(\epsilon*y_{ik})(t),\qquad(\epsilon*y)(t):=\int_{0}^{t}\epsilon(t-s)\,y(s)\,ds.$ (B.2)
Differentiating w.r.t. $Z_{ij}$ gives
$e_{ij}^{Z}(t)=\frac{\partial u_{i}(t)}{\partial Z_{ij}}=(\epsilon*\psi_{ij}^{Z})(t),\qquad\psi_{ij}^{Z}(t):=\frac{\partial y_{ij}(t)}{\partial Z_{ij}}=\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$ (B.3)
B.2 Exponential EPSP kernel: ODE / event-driven update
For $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the convolution representation (B.3) is equivalent to the linear ODE
$\dot{e}_{ij}^{Z}(t)=-\frac{1}{\tau_{m}}e_{ij}^{Z}(t)+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$ (B.4)
Hence $e_{ij}^{Z}(t)$ decays exponentially between presynaptic spikes and exhibits jumps at spike times:
$\begin{cases}\text{between spikes:}\quad&e_{ij}^{Z}(t+\Delta t)=e_{ij}^{Z}(t)\,e^{-\Delta t/\tau_{m}},\\ \text{at }t=t_{j}^{f}:\quad&e_{ij}^{Z}(t^{+})=e_{ij}^{Z}(t^{-})+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}.\end{cases}$ (B.5)
B.3 STD dynamics and parameter sensitivities
Under the STD-only assumption of the main text, the synaptic efficacy is
$w_{ij}(t)=w_{ij}^{0}\,U_{ij}\,d_{ij}(t),\qquad d_{ij}(t)\in[0,1],$ (B.6)
and the depression variable evolves as
$\dot{d}_{ij}(t)=\frac{1-d_{ij}(t)}{\tau_{d}}-U_{ij}\,d_{ij}(t^{-})\,x_{j}(t).$ (B.7)
Define the parameter sensitivity of $d_{ij}$ by
$s_{ij}^{Z}(t):=\frac{\partial d_{ij}(t)}{\partial Z_{ij}}.$ (B.8)
For $Z_{ij}=w_{ij}^{0}$, $d_{ij}(t)$ does not depend on $w_{ij}^{0}$, hence
$s_{ij}^{w_{0}}(t)\equiv 0.$ (B.9)
For $Z_{ij}=U_{ij}$, differentiating (B.7) yields the distributional ODE
$\dot{s}_{ij}^{U}(t)=-\frac{1}{\tau_{d}}s_{ij}^{U}(t)-\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right]x_{j}(t),$ (B.10)
which corresponds to the event-driven update
$\begin{cases}\text{between spikes:}\quad&s_{ij}^{U}(t+\Delta t)=s_{ij}^{U}(t)\,e^{-\Delta t/\tau_{d}},\\ \text{at }t=t_{j}^{f}:\quad&s_{ij}^{U}(t^{+})=(1-U_{ij})\,s_{ij}^{U}(t^{-})-d_{ij}(t^{-}).\end{cases}$ (B.11)
Finally, the parameter derivatives of the synaptic efficacy (evaluated at spike arrival times) are
$\frac{\partial w_{ij}(t^{-})}{\partial w_{ij}^{0}}$ $=U_{ij}\,d_{ij}(t^{-}),$ (B.12)
$\frac{\partial w_{ij}(t^{-})}{\partial U_{ij}}$ $=w_{ij}^{0}\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right].$ (B.13)
Substituting (B.12) into (B.4) yields an explicit online computation of $e_{ij}^{Z}(t)$ from the simulated presynaptic spikes and STP state variables.
Appendix C Weak-coupling reduction: factorization and closed STD sensitivity dynamics
This appendix collects derivations that are specific to the weak-coupling baseline ($w=0$): (i) why the score term vanishes at leading order, (ii) how the Poisson compensation yields a factorized form, and (iii) how the same weak-coupling closure implies closed dynamics for the mean synaptic efficacy and the normalized sensitivity functions $f^{w_{0}}$ and $f^{U}$ used in the main text.
C.1 Why the score term vanishes at $w=0$
At $w_{ij}^{0}\equiv 0$, the membrane potentials satisfy $u_{i}(t)=h_{i}(t,\theta)$ deterministically, so $\rho_{i}(t)=\nu_{i}^{0}(t):=g(h_{i}(t,\theta))$ is non-random. Consequently, the Fisher-information functional $\mathcal{J}[X]$ becomes deterministic (i.e., independent of $X$), and the score term in (8) vanishes:
$\begin{split}\left\langle\mathcal{J}[X]\,\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}=\mathcal{J}_{0}\,\left\langle\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}\\ =\mathcal{J}_{0}\,\frac{\partial}{\partial Z_{ij}}\int P(X(T)\mid Z)\,dX=0.\end{split}$ (C.1)
C.2 Compensation formula at $w=0$ and the factorized expectation
At $w=0$, each neuron fires as an inhomogeneous Poisson process with deterministic intensity $\nu_{i}^{0}(t)$. Let $\phi(t)$ be any predictable process (measurable w.r.t. the past spike history up to $t^{-}$). Then the Doob–Meyer decomposition implies the compensation formula
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\left\langle\int_{0}^{T}\phi(t)\,\nu_{j}^{0}(t)\,dt\right\rangle_{w=0}.$ (C.2)
Because $\nu_{j}^{0}(t)$ is deterministic at $w=0$, it can be taken outside the expectation, yielding
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\int_{0}^{T}\nu_{j}^{0}(t)\,\left\langle\phi(t)\right\rangle_{w=0}\,dt.$ (C.3)
Applying this identity to the eligibility representation $e_{ij}^{Z}(t)=\int_{0}^{t}\epsilon(t-t^{\prime})\,\partial_{Z_{ij}}w_{ij}(t^{\prime-})\,dN_{j}(t^{\prime})$ yields the factorized form used in (15):
$\left\langle e_{ij}^{Z}(t)\right\rangle_{w=0}=\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\nu_{j}^{0}(t^{\prime})\,\frac{\partial}{\partial Z_{ij}}\left\langle w_{ij}(t^{\prime})\right\rangle_{w=0}.$ (C.4)
Substituting this into the pathwise term (10) at $w=0$ gives (15).
C.3 Closed dynamics of the mean efficacy and normalized sensitivities
Here we derive the closed dynamics of the normalized sensitivity functions used in the main text. This derivation is conceptually part of the weak-coupling reduction: it relies only on the baseline Poisson statistics (as above) and does not introduce any additional approximation beyond $w\to 0$.
For notational simplicity, we omit synaptic indices and write $w(t)$, $w^{0}$, $U$, and $\nu^{0}(t)$.
C.3.1 Mean dynamics under Poisson spiking
Under STD-only Tsodyks–Markram dynamics, the effective synaptic strength at presynaptic spike times can be written as
$w(t)=w^{0}Ud(t),$ (C.5)
with the equivalent event-driven representation
$\dot{w}(t)=\frac{w^{0}U-w(t)}{\tau_{d}}-Uw(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr),$ (C.6)
where $t^{\text{spike}}$ denotes presynaptic spike times.
Define the conditional mean
$m(t):=\bigl\langle w(t)\bigr\rangle_{X(t)}.$ (C.7)
Under weak coupling, the presynaptic spike train is unaffected by $w^{0}$ and $U$ and can be treated as an inhomogeneous Poisson process with deterministic rate $\nu^{0}(t)$. Then the standard identity holds:
$\bigl\langle w(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr)\bigr\rangle_{X(t)}=\nu^{0}(t)\,\bigl\langle w(t)\bigr\rangle_{X(t)}=\nu^{0}(t)\,m(t).$ (C.8)
Taking the ensemble average yields the closed ODE
$\dot{m}(t)=\frac{w^{0}U-m(t)}{\tau_{d}}-U\nu^{0}(t)\,m(t).$ (C.9)
C.3.2 Parameter sensitivities and normalized sensitivity functions
Differentiate (C.9) with respect to $w^{0}$ (noting $\nu^{0}(t)$ and $U$ do not depend on $w^{0}$):
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial w^{0}}$ $=\frac{U}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial w^{0}}.$ (C.10)
Similarly, differentiation with respect to $U$ gives
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial U}$ $=\frac{w^{0}}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial U}-\nu^{0}(t)\,m(t).$ (C.11)
Because (C.9) is linear in $m(t)$ and proportional to $w^{0}$ (given our initial conditions), the solution satisfies $m(t)=w^{0}\,y(t;U)$ and hence
$\frac{\partial m(t)}{\partial w^{0}}=\frac{m(t)}{w^{0}}.$ (C.12)
Now define the normalized sensitivity functions (main text Eq. (16)):
$f^{w_{0}}(t)$ $:=\frac{1}{U}\,\frac{\partial m(t)}{\partial w^{0}},$ $f^{U}(t)$ $:=\frac{1}{w^{0}}\,\frac{\partial m(t)}{\partial U}.$ (C.13)
Using (C.10) yields
$\dot{f}^{w_{0}}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{w_{0}}(t),$ (C.14)
and using (C.11) together with (C.12) yields
$\dot{f}^{U}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{U}(t)-\nu^{0}(t)\,Uf^{w_{0}}(t).$ (C.15)
These are exactly the dynamics stated in the main text (Eq. (17), with $\nu^{0}(t)$ denoted there as $\nu_{j}^{0}(t)$).
Appendix D Analysis of STD sensitivity functions and frequency-domain response
This appendix analyzes the properties of the sensitivity functions $f^{w_{0}}(t)$ and $f^{U}(t)$ and the effective presynaptic learning term $C^{Z}(t)=f^{Z}(t)\nu(t)$, starting from the closed weak-coupling dynamics derived in Appendix C.3 (equivalently, main text Eq. (17)). For notational simplicity, we omit synaptic indices.
D.1 Response to constant input
For a constant presynaptic firing rate $\nu$, the system of differential equations (17) yields the following steady-state solutions:
$f^{w_{0}}_{*}=\frac{1}{1+\tau_{d}\nu U},\quad f^{U}_{*}=\frac{1}{(1+\tau_{d}\nu U)^{2}}.$ (D.1)
The transient responses from arbitrary initial conditions are given by:
$f^{w_{0}}(t)$ $=f^{w_{0}}_{*}+\left[f^{w_{0}}(0)-f^{w_{0}}_{*}\right]e^{-\kappa t}$ (D.2)
$f^{U}(t)$ $=f^{U}_{*}+\bigg[\left(f^{U}(0)-f^{U}_{*}\right)-\nu U\left(f^{w_{0}}(0)-f^{w_{0}}_{*}\right)t\bigg]e^{-\kappa t},$ (D.3)
where $\kappa=\tau_{d}^{-1}+\nu U$ represents the effective decay rate.
D.2 Linear response to sinusoidal modulation
We analyze the response of the sensitivity functions and the resulting presynaptic learning term to a weak sinusoidal modulation of the presynaptic firing rate. We consider a rate modulation of the form
[…]
where $\nu_{0}$ is the baseline rate, $\hat{\nu}$ is the small amplitude of modulation ($|\hat{\nu}|\ll\nu_{0}$), and $\omega$ is the angular frequency.
D.2.1 Linearization of the dynamics.
We decompose the sensitivity functions into their steady-state values and small time-dependent fluctuations:
$f^{w_{0}}(t)=f^{w_{0}}_{*}+\delta f^{w_{0}}(t),\qquad f^{U}(t)=f^{U}_{*}+\delta f^{U}(t).$ (D.5)
Substituting these into Eq. (17) and retaining only first-order terms in $\delta f$ and $\delta\nu$, we obtain the linearized system:
$\dot{\delta f}^{w_{0}}(t)$ $=-\kappa\delta f^{w_{0}}(t)-Uf^{w_{0}}_{*}\delta\nu(t),$ (D.6)
$\dot{\delta f}^{U}(t)$ $=-\kappa\delta f^{U}(t)-\nu_{0}U\delta f^{w_{0}}(t)-U(f^{U}_{*}+f^{w_{0}}_{*})\delta\nu(t),$ (D.7)
where $\kappa=\tau_{d}^{-1}+\nu_{0}U$ is the effective decay rate derived in the previous subsection.
[…]
$H_{f}^{w_{0}}(\omega)$ $=-\frac{Uf^{w_{0}}_{*}}{\kappa+i\omega},$ (D.8)
$H_{f}^{U}(\omega)$ $=-\frac{U(f^{U}_{*}+f^{w_{0}}_{*})}{\kappa+i\omega}+\frac{\nu_{0}U^{2}f^{w_{0}}_{*}}{(\kappa+i\omega)^{2}}.$ (D.9)
These equations describe low-pass filtering characteristics, where the sensitivity functions effectively integrate the inverted firing rate fluctuations.
For $f^{w_{0}}$, the amplitude gain and phase lag relative to the input modulation can be explicitly derived as:
$|H_{f}^{w_{0}}(\omega)|$ $=\frac{Uf^{w_{0}}_{*}}{\sqrt{\kappa^{2}+\omega^{2}}},$ (D.10)
$\phi_{f}^{w_{0}}(\omega)$ $=\pi-\tan^{-1}\left(\frac{\omega}{\kappa}\right).$ (D.11)
From these expressions, it is evident that the phase lag $\phi_{f}^{w_{0}}(\omega)$ decreases monotonically from $\pi$ to $\pi/2$ as the frequency $\omega$ increases from 0 to $\infty$ (Figure S1).
For $f^{U}$, although the explicit decomposition is algebraically more complex, the phase lag similarly exhibits a monotonic decrease from $\pi$ to $\pi/2$. Crucially, however, due to the contribution of the second-order pole term, $f^{U}$ responds more sluggishly than $f^{w_{0}}$. Consequently, in the intermediate frequency range where $\omega\sim\kappa$, the phase lag of $f^{U}$ is consistently larger than that of $f^{w_{0}}$ (Figure S1).
Figure S1: Frequency-domain response of the raw sensitivity functions $f^{w_{0}}$ and $f^{U}$. Amplitude gain (upper) and phase lag (lower) of $f$ (not the product $C=f\nu$). These functions behave as low-pass filters with different effective orders. Parameters: $\tau_{d}=0.5$~s, $\nu_{0}=10$~Hz, $U=0.15$.
D.2.2 Response of the effective presynaptic term.
The synaptic learning rule (19) depends on the effective presynaptic term $C^{Z}(t)=f^{Z}(t)\nu(t)$. Its linearization yields:
$\delta C^{Z}(t)\approx f^{Z}_{*}\delta\nu(t)+\nu_{0}\delta f^{Z}(t).$ (D.12)
The corresponding frequency response function $H_{C}^{Z}(\omega)=\hat{C}^{Z}/\hat{\nu}$ is therefore given by the superposition of the direct rate modulation and the filtered sensitivity dynamics:
$H_{C}^{Z}(\omega)=f^{Z}_{*}+\nu_{0}H_{f}^{Z}(\omega).$ (D.13)
D.2.3 Response characteristic for $Z=w_{0}$.
For the baseline weight parameter ($Z=w_{0}$), substituting $H_{f}^{w_{0}}(\omega)$ yields:
$H_{C}^{w_{0}}(\omega)$ $=f^{w_{0}}_{*}\left(1-\frac{\nu_{0}U}{\kappa+i\omega}\right)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}+i\omega}{\kappa+i\omega}.$ (D.14)
This transfer function represents a lead–lag compensator. Since $\kappa>\tau_{d}^{-1}$, the phase is positive (leading) for all $\omega>0$. Writing the phase as the difference between the arguments of the zero and the pole,
$\phi(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right),$ (D.15)
its stationary points satisfy
$\frac{\mathrm{d}\phi}{\mathrm{d}\omega}=\frac{\tau_{d}^{-1}}{\tau_{d}^{-2}+\omega^{2}}-\frac{\kappa}{\kappa^{2}+\omega^{2}}=0.$ (D.16)
Solving this equation gives
$\omega^{2}=\tau_{d}^{-1}\kappa,$ (D.17)
and hence the phase lead is maximized at the geometric mean of the pole and zero frequencies,
$\omega_{*,w_{0}}=\sqrt{\tau_{d}^{-1}\kappa}.$ (D.18)
Because $\phi(\omega)\to 0$ as $\omega\to 0$ and as $\omega\to\infty$, and $\phi(\omega)>0$ for intermediate frequencies, this stationary point corresponds to the global maximum of the phase lead.
It is also instructive to examine the limiting cases of $H_{C}^{w_{0}}(\omega)$. In the quasi-static limit $\omega\to 0$, we obtain
$H_{C}^{w_{0}}(0)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}}{\kappa}=(\kappa\tau_{d})^{-2}.$ (D.19)
Using $\kappa\tau_{d}=1+\tau_{d}\nu_{0}U$, this low-frequency gain can be written as
$H_{C}^{w_{0}}(0)=\left.\frac{\mathrm{d}}{\mathrm{d}\nu}\bigl[\nu f_{*}^{w_{0}}(\nu)\bigr]\right|_{\nu=\nu_{0}}=f_{*}^{w_{0}}(\nu_{0})+\nu_{0}\frac{\partial f_{*}^{w_{0}}}{\partial\nu_{0}},$ (D.20)
that is, it coincides with the derivative of the steady-state effective presynaptic term $C_{*}^{w_{0}}(\nu)=\nu f_{*}^{w_{0}}(\nu)$ with respect to $\nu$, evaluated at $\nu_{0}$. Thus, when the input varies sufficiently slowly, the response of $C^{w_{0}}(t)$ is consistent with a quasi-static approximation in which the system adiabatically tracks the steady-state relation between $\nu$ and $C^{w_{0}}$.
In the opposite limit $\omega\to\infty$, the transfer function reduces to
$\lim_{\omega\to\infty}H_{C}^{w_{0}}(\omega)=\frac{1}{\tau_{d}\kappa}=f^{w_{0}}_{*}.$ (D.21)
Here the oscillatory modulation is too fast for the sensitivity dynamics to follow, so the fluctuation $\delta f^{w_{0}}(t)$ is effectively averaged out. Consequently, only the instantaneous modulation of the firing rate contributes and the gain of the effective presynaptic term is given by the constant steady-state factor $f^{w_{0}}_{*}$.
D.2.4 Response characteristic for $Z=U$.
For the release probability parameter ($Z=U$), the derivation is more involved due to the coupled dynamics. Substituting $H_{f}^{U}(\omega)$ and simplifying the algebraic terms, we obtain:
$H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}+i\omega}{(\kappa+i\omega)^{2}}\left(-\tau_{d}\kappa^{2}+2\kappa+i\omega\right).$ (D.22)
To analyze the phase, it is convenient to introduce the dimensionless parameters
$r:=\kappa\tau_{d},\qquad x:=\frac{\omega}{\kappa}.$ (D.23)
Using $\tau_{d}^{-1}=\kappa/r$ and $\tau_{d}\kappa^{2}+2\kappa=\kappa(2-r)$, the transfer function can be rewritten as
$H_{C}^{U}(\omega)=\frac{1}{r^{2}}\,\frac{\bigl(\frac{1}{r}+ix\bigr)\bigl((2-r)+ix\bigr)}{(1+ix)^{2}},$ (D.24)
where the positive prefactor $1/r^{2}$ does not affect the phase. Thus, the phase response $\phi_{U}(\omega)=\arg H_{C}^{U}(\omega)$ can be expressed as
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x).$ (D.25)
The stationary points of the phase satisfy
$\frac{\mathrm{d}\phi_{U}}{\mathrm{d}x}=\frac{r}{1+r^{2}x^{2}}+\frac{2-r}{(2-r)^{2}+x^{2}}-\frac{2}{1+x^{2}}=0.$ (D.26)
Solving this equation for $x^{2}$ yields a closed-form expression for the frequency at which the phase is extremal. Substituting $x_{*}=\omega_{*,U}/\kappa$ and rearranging in terms of $r=\kappa\tau_{d}$, we obtain
$\omega_{*,U}^{2}=\kappa^{2}\,\frac{-r(r-2)(r-1)+\sqrt{r(r-2)\bigl\{(r-1)^{4}-4\bigr\}}}{r(r+1)}.$ (D.27)
A real solution for $\omega_{*,U}$ exists only when the expression under the square root is positive and the right-hand side is non-negative. Simple, but somewhat tedious calculation reveals the extremum $\omega_{*,U}$ exists for $1<r<2$ and $r>3$.
The limiting behavior of $H_{C}^{U}(\omega)$ is similar to that of $H_{C}^{w_{0}}(\omega)$. In the quasi-static limit $\omega\to 0$, we find
$\lim_{\omega\to 0}H_{C}^{U}(\omega)$ $=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}}{\kappa^{2}}\,\bigl(-\tau_{d}\kappa^{2}+2\kappa\bigr)$ (D.28)
$=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}.$ (D.29)
Recalling that at steady state the sensitivity for $U$ is
$f_{*}^{U}(\nu)=\frac{1}{\bigl(1+\tau_{d}\nu U\bigr)^{2}},$ (D.30)
the corresponding effective presynaptic term reads
$C_{*}^{U}(\nu)=\nu f_{*}^{U}(\nu)=\frac{\nu}{\bigl(1+\tau_{d}\nu U\bigr)^{2}}.$ (D.31)
A straightforward calculation then shows
$\left.\frac{\mathrm{d}C_{*}^{U}}{\mathrm{d}\nu}\right|_{\nu=\nu_{0}}=\frac{1-\tau_{d}\nu_{0}U}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{3}}=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}=\lim_{\omega\to 0}H_{C}^{U}(\omega),$ (D.32)
confirming that quasi-static approximation is valid again.
In the opposite limit of fast modulation, $\omega\to\infty$, we obtain
$\lim_{\omega\to\infty}H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}=\frac{1}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{2}}=f_{*}^{U},$ (D.33)
again exhibiting that too rapid fluctuations are effectively averaged out and the gain reduces to the constant factor $f_{*}^{U}$, i.e., the steady-state sensitivity of the synapse to changes in $U$.
D.2.5 Dependence on oscillation frequency and STP operating point.
The small-signal responses derived above depend on the biophysical parameters only through the dimensionless operating point
$r:=\tau_{d}\kappa=1+\tau_{d}\nu_{0}U$ (D.34)
and the dimensionless frequency
$x:=\frac{\omega}{\kappa}.$ (D.35)
For weak sinusoidal input $\nu(t)=\nu_{0}+\hat{\nu}\cos(\omega t)$ with $\hat{\nu}\ll\nu_{0}$, the gain and phase of both the sensitivity $f^{Z}(t)$ and the effective presynaptic term $C^{Z}(t)=f^{Z}(t)\nu(t)$ are completely determined by $(r,x)$; different combinations of $(\nu_{0},U,\tau_{d})$ that yield the same $r$ produce identical frequency responses.
For the baseline-weight component ($Z=w_{0}$), the qualitative behavior is independent of $r$. As shown analytically above, the phase
$\phi_{w_{0}}(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right)$ (D.36)
vanishes in both the quasi-static ($\omega\to 0$) and fast-modulation ($\omega\to\infty$) limits, and exhibits a single positive maximum at $\omega_{*,w_{0}}=\sqrt{\tau_{d}^{-1}\kappa}$. Thus, $C^{w_{0}}(t)$ always shows a band-limited phase lead relative to $\nu(t)$, with the strongest lead at intermediate frequencies where the zero at $\omega\sim\tau_{d}^{-1}$ and the pole at $\omega\sim\kappa$ interact most strongly. The gain $|H_{C}^{w_{0}}(\omega)|$ increases monotonically from the quasi-static value $(\kappa\tau_{d})^{-2}$ to the high-frequency limit $f_{*}^{w_{0}}=1/(\tau_{d}\kappa)$, so slow modulations are attenuated, whereas faster modulations are relatively enhanced.
By contrast, for the release-probability component ($Z=U$) the shape of the phase response depends qualitatively on $r$. The analytic expression
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x)$ (D.37)
yields three distinct regimes (Figs. S2S4). When $1<r<2$, the phase behaves similarly to the $w_{0}$ case: $\phi_{U}(\omega)$ is zero at $\omega\to 0$ and $\omega\to\infty$ and exhibits a single positive maximum at an intermediate frequency $\omega_{*,U}$. The gain $|H_{C}^{U}(\omega)|$ increases monotonically with $\omega$, as in the $w_{0}$ case. For $2<r<3$, the phase is approximately $\phi_{U}(\omega)\approx\pi$ at very low frequencies and decreases monotonically toward 0 as $\omega$ increases; thus $C^{U}(t)$ is nearly in anti-phase with the slow input and becomes in-phase for rapid modulations. The gain remains monotonically increasing. For $r>3$, the phase curve crosses the branch cut at $\pm\pi$; when the phase is plotted as a continuous (unwrapped) branch, $\phi_{U}(\omega)$ again exhibits an extremum at $\omega_{*,U}$, but now interpolates from $\phi_{U}\approx-\pi$ at very low frequencies to $\phi_{U}\approx 0$ at high frequencies. In physiologically relevant ranges of $r$, this means that $C^{U}(t)$ can be almost perfectly anti-phase with slow inputs, show a pronounced positive phase lead at intermediate frequencies, and become nearly synchronous with $\nu(t)$ at high frequencies. In the same regime, the gain $|H_{C}^{U}(\omega)|$ exhibits a clear maximum at intermediate frequencies, indicating a band-pass-like sensitivity of $C^{U}(t)$ to the onset of rate changes.
Comparing the two components, the effective presynaptic term driven by $w_{0}$ always has the larger gain, $|H_{C}^{w_{0}}(\omega)|>|H_{C}^{U}(\omega)|$, whereas the phase lead (after unwrapping the branch cut at $\pm\pi$) is systematically larger for $U$ than for $w_{0}$ over the relevant frequency range. Finally, increasing $r$ reduces the gain at all frequencies for both parameters, reflecting the stronger overall depression at higher firing rates and release probabilities, while at the same time it enhances the magnitude of the phase shift: depression is engaged earlier within each cycle, so the peak of $C^{Z}(t)$ shifts toward the rising flank of $\nu(t)$.
Figure S2: Frequency response for $1<r<2$. Frequency-domain response of the sensitivity functions $f^{w_{0}}(t)$ and $f^{U}(t)$ (left column) and the effective presynaptic terms $C^{w_{0}}(t)$ and $C^{U}(t)$ (right column) for an operating point with $1<r=\tau_{d}\kappa<2$ (here $r=1.75$, corresponding to $\nu_{0}=10$ Hz, $\tau_{d}=0.5$ s, and $U=0.15$). The presynaptic rate is weakly modulated as $\nu(t)=\nu_{0}+\hat{\nu}\cos(\omega t)$ with $\hat{\nu}\ll\nu_{0}$. Top row: gain $|H_{f}^{Z}(\omega)|$ and $|H_{C}^{Z}(\omega)|$. Bottom row: phase $\phi_{f}^{Z}(\omega)$ and $\phi_{C}^{Z}(\omega)$ (arguments of the corresponding transfer functions), where positive values indicate a phase lead of the output relative to $\nu(t)$. Blue curves: $Z=w_{0}$; orange curves: $Z=U$. Solid lines show the analytical linear-response predictions, and dots show results from numerical simulations of the full nonlinear dynamics. In this regime ($1<r<2$), both $C^{w_{0}}$ and $C^{U}$ exhibit a single positive phase maximum at intermediate frequencies and monotonic increase of the gain with $\omega$.
Figure S3: Frequency response for $2<r<3$. Same layout and conventions as in Fig. S2, but for an operating point with $2<r=\tau_{d}\kappa<3$ (here $r=2.5$; the precise values of $\nu_{0}$ and $U$ are chosen to satisfy this relation). For $Z=w_{0}$, the gain and phase closely resemble those in the $1<r<2$ regime. For $Z=U$, however, the phase of both $f^{U}$ and $C^{U}$ is close to $\pi$ (almost anti-phase) at very low frequencies and decreases monotonically toward 0 as $\omega$ increases, while the gain continues to increase monotonically. Thus, in this regime the $U$-dependent contribution to the effective presynaptic term inverts very slow input modulations but becomes in-phase with the input at high frequencies.
Figure S4: Frequency response for $r>3$. Same layout and conventions as in Fig. S2, but for an operating point with $r=\tau_{d}\kappa>3$ (Here $r=6$ and other parameters are chosen accordingly). For $Z=w_{0}$, the gain and phase again follow the generic pattern of a single positive phase maximum and monotonic increase of the gain. For $Z=U$, the phase curves of $f^{U}$ and $C^{U}$ cross the branch cut at $\pm\pi$ in the principal-value representation, resulting in an apparent discontinuity. When the phase is unwrapped to follow a continuous branch, it interpolates from $\phi\approx-\pi$ at very low frequencies, through a pronounced positive phase lead at intermediate frequencies (peaked at $\omega_{*,U}$), to $\phi\approx 0$ at high frequencies. In the same regime, the gain $|H_{C}^{U}(\omega)|$ exhibits a clear maximum at intermediate frequencies, indicating that the $U$-dependent contribution of the effective presynaptic term is band-pass-like and selectively emphasizes the rising phase of the input.
Figure S5: Comparison across different operating points $r$. Frequency responses of the sensitivity functions and effective presynaptic terms for several values of the operating point $r=\tau_{d}\kappa$ (different curves, indicated in the legend). Upper panels: gain $|H_{f}^{Z}(\omega)|$ and $|H_{C}^{Z}(\omega)|$ for $Z\in\{w_{0},U\}$. Lower panels: corresponding phases $\phi_{f}^{Z}(\omega)$ and $\phi_{C}^{Z}(\omega)$. Left column: $Z=w_{0}$. Right column: $Z=U$. For both parameters, increasing $r$ decreases the gain at all frequencies, reflecting stronger overall short-term depression at higher firing rates and release probabilities. At the same time, the magnitude of the phase shift (relative lead or lag) increases over a broad range of intermediate frequencies, indicating that depression is recruited earlier within each cycle when the synapse operates in a more strongly depressed regime.
Finally, to verify the robustness of these conclusions, we also examined the case of finite-amplitude modulation, $\hat{\nu}=\nu_{0}$. Although the numerical results then deviate quantitatively from the linear-response predictions, the overall frequency dependence of both gain and phase is essentially unchanged: the relative ordering of $|H_{C}^{w_{0}}|$ and $|H_{C}^{U}|$, the presence or absence of an intermediate-frequency maximum, and the characteristic phase lead/lag patterns across $r$ remain intact (Fig. S6).
Figure S6: Effect of finite-amplitude modulation. Frequency responses of the sensitivity and effective presynaptic terms for a larger modulation amplitude, $\nu(t)=\nu_{0}+\hat{\nu}\cos(\omega t)$ with $\hat{\nu}=\nu_{0}=10$ Hz. The parameters match those used in the main-text simulations (Fig. 2) and in Figs. S2S5. Panel layout and color conventions are identical to the previous figures: blue for $Z=w_{0}$, orange for $Z=U$, solid lines for analytical predictions based on linear response, and dots for numerical simulations. Although the larger modulation amplitude leads to quantitative deviations from the linear-response curves, the qualitative behavior of both gain and phase is unchanged: the relative ordering of $|H_{C}^{w_{0}}|$ and $|H_{C}^{U}|$, the presence or absence of an intermediate-frequency maximum in the gain, and the characteristic phase-lead patterns across different operating points $r$ remain essentially the same.
Appendix E Supplementary Figures
Figure S7: Optimal $U(z)$ under different constraints. Optimization results for $U(\zeta)$ under various restrictions. The overall trend remains unchanged by the type of restrictions. A. different sparsity level. B. different mean $\frac{1}{2\pi}\int_{-\pi}^{\pi}d\zeta U(\zeta)$. C. different $\max U(z)$. Other parameters are identical to the figures in the subsection 3.2.
Figure S8: Optimal $U(z)$ under different input amplitudes. Optimization results for $U(\zeta)$ with different amplitudes of input $h(z,t)=A[\cos\omega t-\cos\theta_{c}]$. Stronger input results in more skewed connectivity. Blue: $A=1.0$. Orange: $A=2.0$. Green: $A=3.0$. Other parameters are identical to the figures in the subsection 3.2.
[…]